OpenStudy (anonymous):

how do you graph y=2x^2 on the same coordinate as the parent function?

OpenStudy (jotopia34):

The equation y=2x^2, has a parent function of a parabola y=x^2, so...

OpenStudy (jotopia34):

you would first graph y=x^2, and then do y=2x^2 on the same graph. You will find that the 2x^2 is wider or narrower, I have to plug in some points